OpenStudy (anonymous):

4/5b-2=-9/10b+3 Solve Linear Equation and formula. I BEEN STRUGGLING PLEASE HELP

OpenStudy (tkhunny):

Have you considered multiplying by 10 to make your life simpler?

OpenStudy (tkhunny):

Multiply EVERYTHING by 10. Please feel free to respond on the public boards.

OpenStudy (anonymous):

40/50b-20=-90/100b-30 Like that?

OpenStudy (tkhunny):

?? No, correctly. Practice: 8*(1/2) = ??

OpenStudy (anonymous):

4

OpenStudy (tkhunny):

Why did you not get 1/2? That's what you did to the equations! 8(1/2) = \(\dfrac{8}{1}\cdot\dfrac{1}{2} = \dfrac{8*1}{1*2} = \dfrac{8}{2} = 4\) Go try that "multiply by 10" again. This time, don't you dare let 10 get into anyone's denominator.

OpenStudy (anonymous):

\[\frac{ 4 }{ 5 }b-2=-\frac{ 9 }{ 10 }b+3\]

OpenStudy (anonymous):

Do i multiply the two like terms?

OpenStudy (tkhunny):

Why are you not just multiplying by 10? \(10\cdot\dfrac{4}{5} = 8\) You do the rest.

OpenStudy (anonymous):

8b-2=-9b+3

OpenStudy (anonymous):

b=1 RIGHT?

OpenStudy (anonymous):

hellooooo

OpenStudy (anonymous):

Is anyone here

OpenStudy (tkhunny):

Try again. Remember your Distributive Property. \(10\cdot\left(\dfrac{4}{5}b - 2\right) = 10\cdot\dfrac{4}{5}b - 10\cdot2\) We ARE getting closer.

OpenStudy (anonymous):

8b-20

OpenStudy (tkhunny):

Super. Now the other side.

OpenStudy (anonymous):

-9b+30

OpenStudy (tkhunny):

8b - 20 = -9b + 30 It is hoped that this is an easier problem to solve than the one with which we started. No?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now what do i do. do i subtract -9 on both sides

OpenStudy (tkhunny):

This method, the removal of fractions, CAN be very useful. You should NOT begin to believe that you should do it all the time. Just keep it in your head as something you MIGHT do if it appears to be to your advantage.

OpenStudy (anonymous):

will the answer be -17/50

OpenStudy (anonymous):

Question:how do you know that you were suppost to multiply by 10

OpenStudy (tkhunny):

There is no -9 to subtract. The term is -9b. Not quite. Try adding 9b to both sides, again and adding 20 to both sides again.

OpenStudy (anonymous):

i get 17 and 50

OpenStudy (anonymous):

17b-50

OpenStudy (tkhunny):

8b - 20 = -9b + 30 Add 8b to both sides 8b - 20 + 9b = -9b + 30 + 9b 8b + 9b - 20 = -9b + 9b + 30 17b - 20 = 0b + 30 17b - 20 = 30 Add 20 to both sides 17b - 20 + 20 = 30 + 20 17b - 0 = 50 17b = 50 Divide by 17 17b/17 = 50/17 (17/17)b = 50/17 1b = 50/17 b = 50/17 For some odd reason, you gathered everything to one side and then discarded the equal sign. Don't do that. If you find your equal sign missing, something went horribly wrong. How do you add these? \(\dfrac{1}{5} + \dfrac{1}{2} + \dfrac{1}{10}\)?